OpenStudy (anonymous):

If the population of a country in 2006 equaled 33.8 million, which represented a 4.5% increase from 2001, then how do I find the population of 2001?

OpenStudy (anonymous):

x * 1.045 = 33.8 => x = 33.8 / 1.045 = 32.3 million